OpenStudy (anonymous):

Admission to the Pie Eating Contest is $1.50 for children and $4 for adults. On a certain day, 2200 people came to watch and $5050 is collected. How many children and adults attended the Contest? Show ALL your work. Let x = number of children Let y = number of adults A. Write 2 equations to model this system: 10 points B. Solve the first equation for x (get the x alone on 1 side): 5 points C. Use Substitution to solve the equations and determine how many adults and how many children attended: 5 points

OpenStudy (anonymous):

WILL FAN & MEDAL

OpenStudy (anonymous):

A. x(1.5)+y(4)=5050 1.5/4= z/100 : 4z=150 150/4=37.5=z .375(5050)/1.5=x=1262.5 1262.5(1.5)+y(4)=5050 1893.75+4(y)=5050 4(y)=5050-1893.75 4(y)=3156.25 y=3156.25/4 y=789.0625 x=1262.5 y=789.0625

OpenStudy (anonymous):

(1262.5*1.5)+(789.0625*4)=5050

OpenStudy (anonymous):

I ruined it, don't listen to my answer, I didn't account for 2200 people, I just picked any value that worked

OpenStudy (anonymous):

x+y=2200 x(1.5)+y(4)=5050 2x(1.5)+2y(4)=7250 3.5x+6y=7250 3.5/6=z/100 350=6z z= percentage of kids I hope this helps, I have to go to work, and I can't do anymore right now

OpenStudy (anonymous):

z*7250=x (100-z)*7250=y